OpenStudy (anonymous):

Solve the systems using substitution. 5x - 4y = 14 x - 2y = -2

OpenStudy (anonymous):

solve the second equation for x, then replce in the first equation

OpenStudy (anonymous):

5x - 4y = 14 x - 2y =2 OR x = 2 + 2y Substitute and solve. 5(2 + 2y) - 4y = 14

OpenStudy (anonymous):

equation 1 : 5x - 4y = 14 equation 2 : x - 2y = -2 -> x= 2y -2 So substitute 2y - 2 in place of x in the first equation and solve for y. 5(2y-2)-4y = 14 10y - 10 - 4y = 14 6y = 14+ 10 y = 24/6 y = 4 Then plug in the y value into equation 2 and solve for x. x = 2y-2 x=2(4)-2 x=6

OpenStudy (anonymous):

Oh, it's - 2. o: